# BÅehová 7, Prague, Czech Republic

### v1

@@ 2 Multi-fluid model
Dynamics of heavy-ion reactions in the multi-fluid model (3-fluid model specifically) is desrcibed by a set of coupled local energy-momentum (non-)conservation equations:
@@
each equation represents evolution of a fluid with a source term describing local energy-momentum exchange with the two other fluids. The indices “p” and “t” stand for projectile and target fluids which represent the original nuclei, and “t” stands for the fireball fluid, which represents meson-dominated medium created in the collision.
The three equations (1) summed together correspond to exact energy-momentum conservation of the entire system:
@@ 3 Results and discussion
Overall, MUFFIN reproduces the shape of the $dN_{\rm ch}/d\eta$ distributions while slightly missing the magnitude. The shape of net proton distribution from NA49 experiment is reproduced well, which indicates a correct amount of baryon stopping in the present version of the friction terms. The transverse momentum distributions are reproduced generally well, with a notable exception of anti-proton ones, where the spectra are flatter than the data.
@@
Acknowledgements: the authors acknowledge support by the Czech Science Foundation under project No. 22-25026S.

### v2

@@ 2 Multi-fluid model
Dynamics of heavy-ion reactions in the multi-fluid model (3-fluid model specifically) is desrcibed by a set of coupled local energy-momentum and baryon charge (non-)conservation equations:
@@
$\partial_{\mu}N^{\mu}_{\mathrm{p,t,f}}(x)$ $=0,$
each equation represents evolution of a fluid with a source term describing local energy-momentum exchange with the two other fluids. The indices “p” and “t” stand for projectile and target fluids which represent the original nuclei, and “f” stands for the fireball fluid, which represents meson-dominated medium created in the collision. Projectile and target fluids are initialised as moving blobs with local rest frame energy $\varepsilon_{0}$ and baryon $n_{0}$ densities equal to those of ordinary nuclear matter, whereas the fireball fluid is empty at the beginning.
The first three equations (1) summed together correspond to exact energy-momentum conservation of the entire system:
@@ 3 Results and discussion
Overall, MUFFIN reproduces the shapes of both the $dN_{\rm ch}/d\eta$ and net proton distributions while slightly missing the magnitudes of both. The shape of net proton distribution from NA49 experiment is reproduced well, which indicates a correct amount of baryon stopping in the present version of the friction terms. The transverse momentum distributions are reproduced generally well, with a notable exception of anti-proton ones, where the spectra are flatter than the data.
@@
Acknowledgements: IK and BT acknowledge support by the Czech Science Foundation under project No. 22-25026S. PH was supported by the program Excellence Initiative–Research University of the University of Wrocław of the Ministry of Education and Science. BT acknowledges support by VEGA 1/0521/22. Computational resources were provided by the e-INFRA CZ project (ID:90254), supported by the Ministry of Education, Youth and Sports of the Czech Republic.